OpenStudy (anonymous):

If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all? Numerical Answers Expected!

OpenStudy (anonymous):

We are given C = $145 for Arianna, so: 20 + 25 + 50(x - 2) = 145 50(x - 2) = 145 - 20 - 25 50x - 100 = 100 50x = 200 x = 200/50 x = 4 Since x = 4, Arianna checked 4 bags.